OpenStudy (anonymous):

In order to ensure the safety of school classrooms the local Fire Marshall does an inspection at Thomas Jefferson High School every month, looking for faulty wiring, overloaded circuits, etc. At TJHS the new Academic Wing has 5 math rooms, 10 science rooms, and 10 English rooms. The science rooms are divided into 8 biology and 2 chemistry rooms. Each month, the Fire Marshall randomly picks one of the rooms in the new wing to inspect each month. Define the following events: S = the event the selected room is a science room B = the event the selected room is a biology room M = the event the s

OpenStudy (anonymous):

In order to ensure the safety of school classrooms the local Fire Marshall does an inspection at Thomas Jefferson High School every month, looking for faulty wiring, overloaded circuits, etc. At TJHS the new Academic Wing has 5 math rooms, 10 science rooms, and 10 English rooms. The science rooms are divided into 8 biology and 2 chemistry rooms. Each month, the Fire Marshall randomly picks one of the rooms in the new wing to inspect each month. Define the following events: S = the event the selected room is a science room B = the event the selected room is a biology room M = the event the selected room is a math room E = the event the selected room is an English room C = the event the selected room is a chemistry room Calculate the probabilities of the events described below: a) P(S) b) P(M or E) c) P(E or B) d) P(S and not C)

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

5 M 10 S (8 B, 2 C) 10 E

ganeshie8 (ganeshie8):

So, in total there are 25 rooms, eh ?

ganeshie8 (ganeshie8):

out of which 10 rooms belong to Science, a) P(S) = 10/25 = 2/5

ganeshie8 (ganeshie8):

see if u can work part b

OpenStudy (anonymous):

for b would it be P(M or E)= (5*10)/25= 2

ganeshie8 (ganeshie8):

Nope, u need to add

ganeshie8 (ganeshie8):

number of math or english rooms = 5+10 = 15 total number of rooms = 25 so, P(M or E) = 15/25 = 3/5

OpenStudy (anonymous):

Ok so C) English or biology room will be = 10+8= 18 divided by the total number of room so it will be 18/25 so P(E or B) = 18/25=.75 is that correct?

ganeshie8 (ganeshie8):

thats \(\large \color{red}{\checkmark}\)

ganeshie8 (ganeshie8):

u made a minor calculation mistake

ganeshie8 (ganeshie8):

corrected below : P(E or B) = 18/25=.72

ganeshie8 (ganeshie8):

okay ?

OpenStudy (anonymous):

oh yeah i just look at the calculator wrong

OpenStudy (anonymous):

but i got .72

ganeshie8 (ganeshie8):

0.72 is right

OpenStudy (anonymous):

Ok . For C) how would i got about it?

ganeshie8 (ganeshie8):

u want to pick a Science room, but dont want it to be Chemistry

ganeshie8 (ganeshie8):

that means, u want to pick Biology room right ?

ganeshie8 (ganeshie8):

so, simply calculate probability for B

ganeshie8 (ganeshie8):

P(B) = ?

OpenStudy (anonymous):

so P(B)= .32

ganeshie8 (ganeshie8):

Correct !

OpenStudy (anonymous):

Awesome thank you so much for the help and walking me through it

ganeshie8 (ganeshie8):

np.. u wlc :)